[Hallc_running] [New Logentry] RC Update

hszumila at jlab.org hszumila at jlab.org
Sat Nov 6 11:15:01 EDT 2021

Logentry Text:
5-pass high current (70uA) CW beam to Hall C was established over night and went well (as Jay put it). Still high number of trips to sort out. Accelerator will be tuning up beam to A and D today before coming back to work on the C beam and pushing to perhaps even high currents. 

If things go well, then the pass change to 3-pass could happen around tomorrow during the day shift allowing for 3-pass physics beam to Hall C on the swing shift on Sunday. 

For now, TOs are still required and the day and later shift crews are still on stand-by for tomorrow. I'll have an update tomorrow morning around 10am if the day shift leader is required. 

In other news, we are struggling with the HMS rotation, throwing interlocks every time we rotate. Amy has been down in the Hall this morning to help us rotate, but hopefully the issue doesn't return when we start physics. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3937942
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211106/e0402ebe/attachment.html>

More information about the Hallc_running mailing list